# Advancing Physics Data Analysis through Machine Learning and Physics-Informed Neural Networks

### v1

@@ 1 Introduction
Navigating the enigmatic realms of particle physics often leads us to confront phenomena that remain shrouded in mystery, such as the puzzling properties of neutrino masses and the elusive essence of dark matter. These scientific enigmas challenge our current frameworks of understanding and compel the pursuit of more sophisticated theories. To unravel these mysteries, it is essential to test a variety of theoretical and phenomenological models against the experimental data. Our study is propelled to help in the phenomenological study of those models, by implementing a machine learning filter. Our approach harnesses the power of machine learning to refine the investigative lens through which we examine these models.
In the initial stage of the physics analysis, we begin by randomly initializing the free parameters of our theoretical model. These parameters undergo a preliminary process where the tree-level masses are computed, a crucial step for establishing the foundation of our simulation environment. Subsequently, these calculated values are inputted into the SPheno [1, 2] code. SPheno plays a pivotal role in our analysis, as it is tasked with a comprehensive calculation suite: it determines two- and three-body decay modes of both supersymmetric particles and Higgs bosons. Beyond decay calculations, SPheno extends its functionality to evaluate production cross-sections for these particles in electron-positron annihilation scenarios. The versatility of SPheno’s output allows for further analytical exploration through integration with other sophisticated software tools. Notably, it can interface with HiggsBounds [3] and HiggsSignals [4], which scrutinize the compatibility of Higgs boson observations with theoretical predictions, and with MadGraph5_aMC@NLO [5], a tool renowned for simulating high-energy collisions that further elucidates the dynamical behaviours of particles under our model. This methodical and integrated approach, leveraging the strengths of multiple computational tools, furnishes a comprehensive examination of the theoretical model, enabling a detailed exploration of its phenomenological implications which is seen in the flowchart in Fig. 1. The flowchart outlined above enables highly accurate calculations, but this precision comes with a significant time cost. To address this efficiency challenge, we propose incorporating a machine learning (ML) filter at the initial stage of variable initialization. This ML filter is designed to provide a preliminary assessment of whether the randomly initialized variables are likely to be experimentally viable. By applying this filter, we can selectively perform the more time-intensive, precise calculations only on those sets of parameters predicted to be within experimental bounds. This approach aims to streamline the computational process, ensuring that resources are focused on scenarios yielding meaningful results. While the detailed exploration of physics scenarios and the underlying calculations form the cornerstone of our work, they extend beyond the primary focus of this project. As such, we do not delve deeply into the specifics of these physics computations within the current study. For readers interested in a more thorough understanding of these aspects, we direct them to Ferreira et al. [6], where these topics are extensively covered. Our emphasis here is on leveraging machine learning techniques to enhance the efficiency and effectiveness of identifying experimentally viable parameter sets, rather than on the intricacies of the physics calculations themselves.
@@ 2 Dataset Description and Preparation
The dataset employed in this study was derived from the comprehensive phenomenological analysis conducted by Ferreira et al. [6, 7], focusing on a Branco-Grimus-Lavoura (BGL) model. This model, a variant within the two Higgs doublet framework augmented by a single scalar singlet, stands out for its rich phenomenology in particle physics. The original dataset encompasses both experimental outcomes and the corresponding set of free parameters instrumental in computing these values.
For our analysis, we engaged in an extensive data-cleaning process. This refinement involved selecting only the vital free parameters and specific physics values (denoted by $\mathbf{P}$) relevant for predictive modelling. Additionally, the dataset was labelled based on a spectrum of experimental observables, categorizing each instance as either experimentally viable or nonviable according to predefined experimental criteria. This preprocessing was imperative for ensuring that our models concentrated on the free parameters, thus enhancing the accuracy and interpretability of the analysis within the framework of the BGL model. Notably, our dataset contains approximately 16% of viable points, indicating an imbalance. We represent the dataset as $\mathbf{X}$, where $\mathbf{X}=\{x_{1},\dots,x_{N}\}$, and N is the number of simulated data points. Correspondingly, a supplementary dataset contains the physics observables, represented as $\mathbf{P}$, such that for each data point $x_{i}$ with $i\in\{1,\dots,N\}$, there exists an associated set of physics observables $\mathbf{p}_{i}=\{p^{i}_{1},p^{i}_{2},p^{i}_{3}\}$, with $i$ ranging from 1 to N.
@@ 3 Methodology
As part of our study, we conducted a comprehensive evaluation of various machine learning algorithms to determine their effectiveness in predicting outcomes based on our dataset. The assessment process began with the preparation of the dataset, followed by the application of standard scaling to normalize the feature space. This preprocessing step is crucial for models that are sensitive to feature magnitude. We implemented a stratified k-fold cross-validation approach with five folds for evaluating the ML algorithms. This method ensures a robust assessment while maintaining the original dataset’s class proportion within each fold, thus addressing potential biases in model performance due to imbalanced class distributions. Our selection of ML models ranged from simpler algorithms, such as Gaussian Naive Bayes and Decision Trees, to more complex ensemble methods including Random Forest, AdaBoost, and XGBoost [8]. Each model was rigorously tested to gain a comprehensive understanding of its performance, focusing on metrics such as accuracy, precision, recall, F1 score, and the area under the ROC curve (ROC AUC). These metrics were selected for their capacity to provide a detailed insight into model performance, particularly in binary classification tasks where balancing true positive and false positive rates is essential. We systematically recorded the evaluation results, noting not only the mean and standard deviation of each metric across the folds but also the time required for each model’s training and evaluation. This method allowed us to make nuanced comparisons between the models, considering both their predictive performance and computational efficiency.
For the deep neural networks segment of our study, we adopted a focused approach to assess the training procedure, primarily utilizing the validation loss as a key indicator of performance. This method involves dividing our dataset into training, validation, and test subsets, with the latter serving to further validate our results and as a proxy to estimate the model’s generalization ability on unseen data. Throughout the training epochs, we closely monitored the validation loss, aiming to minimize it as an objective measure of the model’s prediction error. A distinctive emphasis was placed on exploring the capabilities of Physics-Informed Neural Networks (PINNs) [9] alongside traditional Neural Networks (NNs) to address the unique challenges presented by our physics-based dataset. In subsection 3.1, we delve into the structural details of these models and their operational dynamics.
@@ 3.1 Physics-Informed Neural Networks (PINNs)
By incorporating this additional physics-based loss term, the PINN is explicitly trained not only to achieve accuracy in its primary tasks but also to approximate physical values and make decisions informed by physics. This approach enhances both the interpretability and reliability of the model’s outputs, bridging the gap between complex data analysis and fundamental physical principles.
@@ 4 Results and Discussion
Table 2: Performance Metrics of Top Neural Network Models Under Different Learning Rates. This table showcases the results for neural network configurations that demonstrated the best overall performance in our analysis. Values in bold represent the highest scores achieved in each respective performance category across all tested models, indicating the most effective model settings for each metric."

### v2

@@ 1 Introduction
Particle physics involves complex phenomena, such as neutrino masses and dark matter, which challenge existing theoretical frameworks. To explore these phenomena, it is essential to test various theoretical models against experimental data. This study focuses on enhancing the efficiency of this testing process by integrating machine learning (ML) techniques. Specifically, we introduce an ML filter to streamline the identification of experimentally viable parameter sets, improving computational efficiency and model validation in particle physics simulations.
The initial stage of the physics analysis involves randomly initializing the free parameters of our theoretical model and computing the tree-level masses. These computed values are then input into the SPheno spheno1 ; spheno2 code, which calculates decay modes for supersymmetric particles and Higgs bosons, as well as production cross-sections in electron-positron annihilation scenarios. SPheno also integrates with other tools, such as HiggsBounds higgsBounds and MadGraph5_aMC@NLO MadGraph , to evaluate the compatibility of Higgs boson observations with theoretical predictions and simulate high-energy particle collisions. This integrated approach enables a comprehensive examination of the model, as shown in Fig.1. While the method offers high accuracy, it comes at a significant computational cost. To improve efficiency, we introduce a machine learning (ML) filter at the initialization stage, which preliminarily assesses whether the randomly initialized variables are likely to be experimentally viable. By applying this filter, we can focus computational resources on promising parameter sets, thus optimizing the overall process. Although the detailed physics computations are integral to our work, they are not the primary focus of this study. For a deeper understanding of the physics calculations, we refer readers to Ferreira et al. ferreira2022phenomenology . Our main emphasis is on leveraging machine learning to enhance the efficiency of identifying experimentally viable parameter sets.
@@ 2 Dataset Description and Preparation
The dataset employed in this study was derived from the comprehensive phenomenological analysis conducted by Ferreira et al. ferreira2022phenomenology ; ferreira2023collider , focusing on a Branco-Grimus-Lavoura (BGL) model. This model, a variant within the two Higgs doublet framework augmented by a single scalar singlet, stands out for its rich phenomenology in particle physics. The original dataset encompasses both experimental outcomes and the corresponding set of free parameters instrumental in computing these values.
For our analysis, we conducted an extensive data-cleaning process, in which we retained only a subset of variables relevant to our work from the initial dataset. This refinement involved selecting the essential free parameters and specific physics values (denoted by $\mathbf{P}$) relevant for predictive modeling. The selected points were tested based on electroweak precision observables (or oblique parameters), Higgs observables, and quark flavor violation (QFV) constraints included in $\mathbf{P}$, categorizing each instance as either experimentally viable or nonviable according to predefined experimental criteria. This preprocessing step was crucial to ensure that our models focused on the free parameters, thereby enhancing the accuracy and interpretability of the analysis within the framework of the BGL model. Notably, our dataset contains approximately 16% viable points, indicating an imbalance. We represent the dataset as $\mathbf{X}$, where $\mathbf{X}={x_{1},\dots,x_{N}}$, and $N$ is the number of simulated data points. A supplementary dataset contains the physics observables, represented as $\mathbf{P}$, such that for each data point $x_{i}$ with $i\in{1,\dots,N}$, there is an associated set of physics observables $\mathbf{p}_{i}={p^{i}_{1},p^{i}_{2},p^{i}_{3}}$, with $i$ ranging from 1 to $N$. Further details regarding the restrictions and the physics can be found in Ferreira et al. ferreira2022phenomenology .
@@ 3 Methodology
As part of our study, we conducted a comprehensive evaluation of various machine learning algorithms to determine their effectiveness in predicting outcomes based on our dataset. The assessment process began with the preparation of the dataset, followed by the application of standard scaling to normalize the feature space. This preprocessing step is crucial for models that are sensitive to feature magnitude. We implemented a stratified k-fold cross-validation approach with five folds for evaluating the ML algorithms. This method ensures a robust assessment while maintaining the original dataset’s class proportion within each fold, thus addressing potential biases in model performance due to imbalanced class distributions. Our selection of ML models ranged from simpler algorithms, such as Gaussian Naive Bayes and Decision Trees, to more complex ensemble methods including Random Forest, AdaBoost, and XGBoost XGBoost . Each model was rigorously tested to gain a comprehensive understanding of its performance, focusing on metrics such as accuracy, precision, recall, F1 score, and the area under the ROC curve (ROC AUC). These metrics were selected for their capacity to provide a detailed insight into model performance, particularly in the classification ”valid point” or ”no-valid point.” tasks where the inherent imbalance in our dataset needs to be addressed, as balancing true positive and false positive rates is essential. Creating synthetic data would introduce bias, and therefore, it was not employed in our study. We systematically recorded the evaluation results, noting not only the mean and standard deviation of each metric across the folds but also the time required for each model’s training and evaluation. This method allowed us to make nuanced comparisons between the models, considering both their predictive performance and computational efficiency.The hyperparameters used for training each of these machine learning models are summarized in Table 1.
For the deep neural networks segment of our study, we adopted a focused approach to assess the training procedure, primarily utilizing the validation loss as a key indicator of performance. This method involves dividing our dataset into training, validation, and test subsets, with the latter serving to further validate our results and as a proxy to estimate the model’s generalization ability on unseen data. Throughout the training epochs, we closely monitored the validation loss, aiming to minimize it as an objective measure of the model’s prediction error. A distinctive emphasis was placed on exploring the capabilities of Physics-Informed Neural Networks (PINNs) pinnRaissi alongside traditional Neural Networks (NNs) to address the unique challenges presented by our physics-based dataset. In subsection 3.1, we delve into the structural details of these models and their operational dynamics.
@@ 3.1 Physics-Informed Neural Networks (PINNs)
Here, $\mathcal{L}_{\text{classification}}$ is given by the Binary Cross-Entropy with Logits Loss function, while $\mathcal{L}_{\text{physics}}$ is the mean-squared-error between the outputs of the physics-informed layer and the physical variables as computed in Ferreira et al. ferreira2022phenomenology . By incorporating this additional physics-based loss term, the PINN is explicitly trained not only to achieve accuracy in its primary tasks but also to approximate physical values and make decisions informed by physics. This approach enhances both the interpretability and reliability of the model’s outputs, bridging the gap between complex data analysis and fundamental physical principles.
@@ 4 Results and Discussion
Model Hyperparameters
Nearest Neighbors n_neighbors = 3
Decision Tree max_depth = 5
Random Forest max_depth = 5, n_estimators = 10, max_features = 1
AdaBoost Default hyperparameters
Naive Bayes Default hyperparameters
QDA Default hyperparameters
XGBoost use_label_encoder = False, eval_metric = ’logloss’
Table 1: Hyperparameters used for training machine learning models of Table 2.
@@
Table 3: Performance Metrics of Top Neural Network Models Under Different Learning Rates. This table showcases the results for neural network configurations that demonstrated the best overall performance in our analysis. Values in bold represent the highest scores achieved in each respective performance category across all tested models, indicating the most effective model settings for each metric.”